OpenStudy (anonymous):

help c(x)=95*1/7*7 please

OpenStudy (unklerhaukus):

what is the question here?

OpenStudy (unklerhaukus):

do you just have to simplify the right hand side?

OpenStudy (anonymous):

yeah

OpenStudy (unklerhaukus):

is it like this?\[c(x)=95\times\frac17\times7 \] or \[c(x)=95\times\frac1{7\times7} \]

OpenStudy (anonymous):

like the first one :)

OpenStudy (unklerhaukus):

\[90\times\frac{1}{7}\times7=95\times\frac{7}{7}=95\times1=\]